Jim is 9 years old. He is older than his sister by 6 years. How many years' time would Jim be twice as old as his sister
Answer by ankor@dixie-net.com(22740)

let j = Jim's present age
let s = sister's
:
write an equation for each statement
:
Jim is 9 years old.
j = 9
He is older than his sister by 6 years.
j = s + 6
replace j with 9, from the first statement
9 = s + 6
s = 9 - 6
s = 3 yrs is the sister's age
:
How many years' time would Jim be twice as old as his sister
let y = no. of years for this to be true
then
j + y = 2(s + y)
j + y = 2s + 2y
j = 2s + 2y - y
j = 2s + y
replace j with 9, replace s with 3
9 = 2(3) + y
9 = 6 + y
9 - 6 = y
y = 3 yrs time, Jim will be twice as old as his sister
:
:
See if that checks out in the statement
" In 3 years' time would Jim be twice as old as his sister
9 + 3 = 2(3 + 3)
12 = 2(6)